Engineers concerned about a tower's stability have done extensive studies of its increasing tilt. Measurements of the lean of the tower over time provide much useful information. The following table gives measurements for the years 1975 to 1987. The variable "lean" represents the difference between where a point on the tower would be if the tower were straight and where it actually is. The data are coded as tenths of a millimeter in excess of 2.9 meters, so that the 1975 lean, which was 2.9642 meters, appears in the table as 642. Only the last two digits of the year were entered into the computer.

Year 75 76 77 78 79 80 81 82 83 84 85 86 87
Lean 642 645 656 667 674 689 696 698 714 717 726 743 758

(a) Plot the data. Consider whether or not the trend in lean over time appears to be linear. (Do this on paper. Your instructor may ask you to turn in this graph.)

(b) What is the equation of the least-squares line? (Round your answers to three decimal places.)
y =  +  x

What percent of the variation in lean is explained by this line? (Round your answer to one decimal place.)
%

(c) Give a 99% confidence interval for the average rate of change (tenths of a millimeter per year) of the lean. (Round your answers to two decimal places.)
(  ,  )
Answer #1

a)

Scatterplot 780 760 740 720 700 680 660 640 620 76 78 80 88 74 82 84 X 86 15

data appear to be linear

b)

X Y (x-x̅)² (y-ȳ)² (x-x̅)(y-ȳ)
75 642 36.000 2728.053 313.385
76 645 25.000 2423.669 246.154
77 656 16.000 1461.592 152.923
78 667 9.000 741.515 81.692
79 674 4.00 409.284 40.46
80 689 1.00 27.361 5.23
81 696 0.00 3.130 0.00
82 698 1.0000 14.2071 3.7692
83 714 4.000 390.822 39.538
84 717 9.000 518.438 68.308
85 726 16.000 1009.284 127.077
86 743 25.000 2378.438 243.846
87 758 36.000 4066.515 382.615
ΣX ΣY Σ(x-x̅)² Σ(y-ȳ)² Σ(x-x̅)(y-ȳ)
total sum 1053 9025 182.000 16172.31 1705.000
mean 81.00 694.23 SSxx SSyy SSxy

sample size ,   n =   13          
here, x̅ = Σx / n=   81.00   ,     ȳ = Σy/n =   694.23  
                  
SSxx =    Σ(x-x̅)² =    182.0000          
SSxy=   Σ(x-x̅)(y-ȳ) =   1705.0          
                  
estimated slope , ß1 = SSxy/SSxx =   1705.0   /   182.000   =   9.3681
                  
intercept,   ß0 = y̅-ß1* x̄ =   -64.5879          
                  
so, regression line is   Ŷ =   -64.588 +   9.368 *x

R² =    (Sxy)²/(Sx.Sy) =    0.988

98..8 percent of the variation in lean is explained by this line

c)

confidence interval for slope                  
α=   0.01              
t critical value=   t α/2 =    3.106   [excel function: =t.inv.2t(α/2,df) ]      
estimated std error of slope = Se/√Sxx =    4.26021   /√   182.00   =   0.316
                  
margin of error ,E= t*std error =    3.106   *   0.316   =   0.981
estimated slope , ß^ =    9.3681              
                  
                  
lower confidence limit = estimated slope - margin of error =   9.3681   -   0.981   =   8.39
upper confidence limit=estimated slope + margin of error =   9.3681   +   0.981   =   10.35